OpenStudy (anonymous):

I WILL MEDAL AND FAN. PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Write the linear equation for cab company (x) in slope-intercept form (y=mx+b), point-slope form ( , and standard form (Ax + By = C). Explain what steps you had to take to manipulate the linear equation into each of the three forms. Cab company (x) will charge $3.25 as a flat fee and $3.00 per every additional mile. Cab company (y) will charge $5.75 as a flat fee and $2.75 per every additional mile

OpenStudy (danjs):

make an equation for each cab based on the information for the total charge for the cabs based on miles Total Charge = Flat Fee + (cost per mile)*(#miles)

OpenStudy (anonymous):

i don't know what the mileage gone is.

OpenStudy (danjs):

ill do Cab X , let cost=y, and miles = x, to look familiar Total Charge = Flat Fee + (cost per mile)*(#miles) y = 3.25 + 3.00*x

OpenStudy (danjs):

mileage is a variable used to determine the cost, it can be any value

OpenStudy (danjs):

notice that is already set up in the y=m*x + b for the first one

OpenStudy (anonymous):

thank you soo much!

OpenStudy (danjs):

point slope y - y1 = m*(x- x1) you have slope m already, m=3, just pick a value for x like 1, and solve for y, to get a point (x1,y1)

OpenStudy (danjs):

x=1 then y=3.25 + 3.00*1 = 6.25 y - 6.25 = 3*(x - 1) , rearange the thing to the standard form, Ax+By = C

OpenStudy (danjs):

try out the other cab, ill be on here for a few min